# CARFT: Boosting LLM Reasoning via Contrastive Learning with Annotated Chain-of-Thought-based Reinforced Fine-Tuning

### v1

@@ Abstract
Wenqiao Zhu1, 2, $\dagger$, Ji Liu1, Rongjuncheng Zhang1, Haipang Wu1, Yulun Zhang2 1HiThink Research, 2Shanghai Jiao Tong University
[…]
@@ 1 Introduction
To address the limitations of SFT-based methods, Reinforcement Learning (RL)-based fine-tuning approaches emerge (Luong et al., 2024; Shao et al., 2024; Liu et al., 2025). A prominent and state-of-the-art RL-based fine-tuning approach is ReFT (Luong et al., 2024), which incorporates online RL approach, i.e., Proximal Policy Optimization (PPO) (Schulman et al., 2017), to dynamically sample CoTs at each training step. This mechanism enables ReFT to leverage multiple CoTs, thereby improving the generalization capability of LLMs.
@@ 2 Related Work
Recent years have witnessed widespread application in Natural Language Processing (NLP), particularly in the domains of preference optimization (Stiennon et al., 2020; Rafailov et al., 2023; Gheshlaghi Azar et al., 2024; Zhu et al., 2025) and reasoning (Luong et al., 2024; Shao et al., 2024; Yu et al., 2025; Liu et al., 2025). These methods typically follow a standard three-stage pipeline: (1) SFT, (2) reward modeling, and (3) RL-based optimization. A key distinction among these approaches lies in how the reward signal is obtained. In preference optimization, reward models are learned from human feedback, while in mathematical reasoning tasks, rule-based methods are typically exploited to construct reward signals, as ground-truth answers can be explicitly verified. Within the context of preference optimization, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has emerged as an effective algorithm that avoids the need for explicit reward model training. However, due to its offline nature (Feng et al., 2024), DPO may struggle to explore diverse CoTs (Luong et al., 2024). As a result, on-policy approaches, e.g., GRPO Shao et al. (2024), DAPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025), and ReFT Luong et al. (2024), are generally employed to better explore such diversity in reasoning.
@@ 3.2.3 Masked Contrastive Signal Construction
Given a batch of training samples $\{\mathbf{x}_{i},\mathbf{c}_{i}^{\text{annotated}},\mathbf{y}_{i}\}_{1}^{B}$ with $B$ presenting the batch size, we conduct LLM self-generation to generate a batch of rollout CoTs, i.e., $\{\mathbf{x}_{i},\mathbf{c}_{i}^{\text{rollout}},\mathbf{y}_{i}\}_{1}^{B}$. By employing the CoT embedding module, we could get embeddings of the annotated CoTs $\{\mathbf{e}^{\text{annotated}}_{i}\}_{1}^{B}$ and rollout CoT $\{\mathbf{e}^{\text{rollout}}_{i}\}_{1}^{B}$ exploiting the approach presented in Section 3.2.2. We construct a contrastive feedback with InfoNCE (Chen et al., 2020) to provide the positive contrastive signal as defined in Formula 7.
@@
where $\mathcal{M}_{2}$ represents a binary mask, in which each element takes the value 1 if the corresponding CoT leads to a wrong answer, and 0 otherwise.
@@ 4.1 Experimental Setup
We conduct experiments on two publicly available datasets: SVAMP (Patel et al., 2021) and GSM8K (Cobbe et al., 2021). Table 2 presents the key statistics of SVAMP and GSM8K. For the reasoning process, we leverage the CoT annotations from (Luong et al., 2024), which were generated based on few-shot prompting (Wei et al., 2022; Gao et al., 2023) with GPT-3.5-turbo (OpenAI, 2023). Our experiments are conducted based on two foundation models: CodeLlama-7B (Rozière et al., 2023) and Qwen2.5-7B-Instruct (Team, 2024). We evaluate CARFT in comparison with three baseline approaches: SFT, ReFT (Luong et al., 2024), and Dr.GRPO (Liu et al., 2025). ReFT is a state-of-the-art RL approach for LLM fine-tuning. As an advanced extension of GRPO (Shao et al., 2024), Dr.GRPO demonstrates excellent performance on R1-like (DeepSeek-AI, 2025) tasks. See setup details in Appendix.
@@ 5 Conclusions
CARFT requires additional computational overhead to compute the embeddings of CoTs, to achieve excellent. As a result, it consumes a longer computational time compared to the ReFT and SFT. However, CARFT needs less computational time than Dr.GRPO as CARFT needs less on-policy sampled CoTs. In addition, CARFT is designed to exploit a centralized annotated CoT dataset. The annotated datasets may be stored in multiple data centers or devices, which may hinder the application of CARFT with decentralized data.
In this work, we have restricted the context to fewer than 1024 tokens; we plan to explore improving reasoning in long-context scenarios (Zhu et al., 2024).
@@ A.2 More Experiments
1
Output : Final Policy: $\mathbf{\pi}_{\theta}$
2 for $i\leftarrow 1$ to $T$ do
 $\hat{\mathbf{c}}\sim\pi_{\theta}$   // On-policy CoT sampling
3
[…]
4
5
6  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{1}$
7 for $i\leftarrow 1$ to $U$ do
8
9   end for
10
11 end for
return $\pi_{\theta}$
Algorithm 1 CARFT with Positive Signal
@@ 4  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{2}$
A.3 CARFT with Positive Signal
[…]
A.4 CARFT with Negative Signal

### v2

@@ Abstract
Wenqiao Zhu1, 2, *, Ji Liu1, *, Rongjunchen Zhang1, Haipang Wu1, Yulun Zhang2 1HiThink Research, 2Shanghai Jiao Tong University
**footnotetext: Corresponding authors: zhuwnq@outlook.com and jiliuwork@gmail.com
@@ 1 Introduction
To address the limitations of SFT-based methods, Reinforcement Learning (RL)-based fine-tuning approaches emerge (Luong et al., 2024; Shao et al., 2024; Liu et al., 2025b). A prominent and state-of-the-art RL-based fine-tuning approach is ReFT (Luong et al., 2024), which incorporates online RL approach, i.e., Proximal Policy Optimization (PPO) (Schulman et al., 2017), to dynamically sample CoTs at each training step. This mechanism enables ReFT to leverage multiple CoTs, thereby improving the generalization capability of LLMs.
@@ 2 Related Work
Reinforcement Learning (RL)-based LLM Reasoning.
Recent years have witnessed widespread application in Natural Language Processing (NLP), particularly in the domains of preference optimization (Stiennon et al., 2020; Rafailov et al., 2023; Gheshlaghi Azar et al., 2024; Zhu et al., 2025) and reasoning (Luong et al., 2024; Shao et al., 2024; Yu et al., 2025; Liu et al., 2025b). These methods typically follow a standard three-stage pipeline: (1) SFT, (2) reward modeling, and (3) RL-based optimization. A key distinction among these approaches lies in how the reward signal is obtained. In preference optimization, reward models are learned from human feedback, while in mathematical reasoning tasks, rule-based methods are typically exploited to construct reward signals, as ground-truth answers can be explicitly verified. Within the context of preference optimization, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has emerged as an effective algorithm that avoids the need for explicit reward model training. However, due to its offline nature (Feng et al., 2024), DPO may struggle to explore diverse CoTs (Luong et al., 2024). As a result, on-policy approaches, e.g., GRPO Shao et al. (2024), DAPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025b), and ReFT Luong et al. (2024), are generally employed to better explore such diversity in reasoning.
@@
Contrastive Learning.
Contrastive learning has shown strong effectiveness in diverse fields such as multimodal pretraining (Radford et al., 2021), recommendation systems (Yang et al., 2022), graph embedding (Zhu et al., 2022), and report generation (Zhou and Wang, 2024). Theoretically, the contrastive loss can be decomposed into two components: an alignment term and a uniformity term (Wang and Isola, 2020). The alignment term reduces the distance between embeddings of positive pairs, while the uniformity term encourages embeddings of negative pairs to be uniformly dispersed in the representation space. Inspired by this framework, we propose in this paper to utilize contrastive feedback to guide the online generation of CoT reasoning.
@@ 3.2.3 Masked Contrastive Signal Construction
1
Output : Final Policy: $\mathbf{\pi}_{\theta}$
2 for $i\leftarrow 1$ to $T$ do
 $\hat{\mathbf{c}}\sim\pi_{\theta}$   // On-policy CoT sampling
3
[…]
4
5
6  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{1}$
7 for $i\leftarrow 1$ to $U$ do
8
9   end for
10
11 end for
return $\pi_{\theta}$
Algorithm 1 CARFT with Positive Signal
@@ 6  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{1}$
[…]
@@
where $\mathcal{M}_{2}$ represents a binary mask, in which each element takes the value 1 if the corresponding CoT leads to a wrong answer, and 0 otherwise. See details in Appendix.
@@ 4.1 Experimental Setup
We conduct experiments on two publicly available datasets: SVAMP (Patel et al., 2021) and GSM8K (Cobbe et al., 2021). Table 2 presents the key statistics of SVAMP and GSM8K. For the reasoning process, we leverage the CoT annotations from (Luong et al., 2024), which were generated based on few-shot prompting (Wei et al., 2022; Gao et al., 2023) with GPT-3.5-turbo (OpenAI, 2023). Our experiments are conducted based on two foundation models: CodeLlama-7B (Rozière et al., 2023) and Qwen2.5-7B-Instruct (Team, 2024). We evaluate CARFT in comparison with three baseline approaches: SFT, ReFT (Luong et al., 2024), and Dr.GRPO (Liu et al., 2025b). ReFT is a state-of-the-art RL approach for LLM fine-tuning. As an advanced extension of GRPO (Shao et al., 2024), Dr.GRPO demonstrates excellent performance on R1-like (DeepSeek-AI, 2025) tasks. See setup details in Appendix.
@@ 4.3 Ablation Study
4.4 Complexity
Simpler methods like SFT require significantly less computational overhead as they do not involve the rollout process used in RL-based methods. However, the performance of SFT is lower compared to RL-based approaches. Our proposed CARFT method achieves the highest performance among all the considered methods.
Let $N$ denote the number of parameters in the LLM, and $L$ represent the length of the on-policy sampled CoT. The computational complexity of a single forward pass in ReFT is $O((L+1)\cdot N)$, while that of SFT is $O(N)$. Since Dr.GRPO requires multiple rollouts, let $G$ be the number of rollouts. Accordingly, the computational complexity of Dr.GRPO becomes $O((G\cdot L+1)\cdot N)$. As for CARFT, which incorporates a contrastive signal, let $M$ represent the size of the projector and $d$ denote the hidden dimension of the LLM. The computational complexity of CARFT is then given by $O((L+1)\cdot N+M\cdot d\cdot L)$, where $M\cdot d$ is significantly smaller than $N$.
@@ 4.4 Complexity
The space complexity of SFT is O(N). In contrast, ReFT incurs an additional overhead of O(Ld) due to caching the sequence of on-policy CoT trajectories, resulting in a total space complexity of $O(N+L\cdot d)$. Similarly, Dr.GRPO has a space complexity of $O(N+G\cdot L\cdot d).$ For CARFT, if we denote the projected hidden size as $d_{2}$, its space complexity becomes $O(N+L\cdot d+M+L\cdot d_{2})$. Here, $M$ is significantly smaller than $N$, and $d_{2}$ (the compressed hidden dimension) is much smaller than the original hidden size $d$.
@@ 5 Conclusions
CARFT requires additional computational overhead to compute the embeddings of CoTs, to achieve excellent. As a result, it consumes a longer computational time compared to the ReFT and SFT. However, CARFT needs less computational time than Dr.GRPO as CARFT needs less on-policy sampled CoTs. In addition, CARFT is designed to exploit a centralized annotated CoT dataset. The annotated datasets may be stored in multiple data centers or devices, which may hinder the application of CARFT with decentralized data. In addition, we have restricted the context to fewer than 1024 tokens in this work. We plan to explore improving reasoning in long-context scenarios (Zhu et al., 2024). In addition, while CARFT exploits centralized data, we will investigate decentralized datasets, e.g., federated learning Liu et al. (2024e, d); Jia et al. (2024); Liu et al. (2024a, c); Che et al. (2023); Liu et al. (2023b, 2022); Zhang et al. (2022); Zhou et al. (2022); Jia et al. (2025); Liu et al. (2025a); Chen et al. (2025); Liu et al. (2024b) and distributed machine learning Liu et al. (2023a).
@@ A.2 More Experiments
Model Model Accuracy
Qwen2.5-Instruct-14B SFT 86.7%
Qwen2.5-Instruct-14B CARFT 88.9%
Table 6: Accuracy curves of various methods on SVAMP dataset and Qwen2.5-14B-Instruct backbone.
Larger Model
In addition to the two 7B models, we conducted experiments on Qwen 2.5 14B, as well. The models were fine-tuned for 100 epochs, and the results are summarized in Table 6. These results highlight the effectiveness and strong generalization of CARFT across diverse model scales and datasets.
@@ 4  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{2}$
A.3 CARFT with Negative Signal
@@
A.3.1 Illustration
We provide an example to illustrate how negative signals are constructed. Let
$A=[a_{1},a_{2},\dots,a_{n}]$
denote the annotated CoT that leads to a correct solution. Let
$B=[b_{1},b_{2},\dots,b_{m}]$
represent the on-policy sampled CoT that results in an incorrect solution.
Step 1: Compute Longest Common Subsequence (LCS)
We first compute the LCS of $A$ and $B$, denoted as:
$C=[c_{1},c_{2},\dots,c_{k}]$
Step 2: Extract Sub-sequences
Next, we compute the embeddings of the following components:
    $C$: the common subsequence,
    $A\setminus C$: parts of the correct CoT not in $C$,
    $B\setminus C$: parts of the incorrect CoT not in $C$.
Step 3: Apply Contrastive Signal
We minimize the contrastive loss to generate the contrastive signal through backpropagation. The motivation behind the negative signal is to align the embedding of $A\setminus C$ with that of $C$, while increasing the distance between the embeddings of $B\setminus C$ and $C$, since $B\setminus C$ leads to an incorrect solution. As noted in (Wang and Isola, 2020), contrastive learning has the net effect of pulling positive pairs together while scattering negative examples apart.
Appendix B Asymptotics of $\mathcal{L}_{\text{contrastive}}$
According to Theorem 1 in (Wang and Isola, 2020), for fixed $\tau>0$, as the number of negative samples $M\to\infty$, the constrative loss converges to
$\lim_{M\to\infty}\mathcal{L}_{\text{contrastive}}(f;\tau,M)-\log M$
$=-\frac{1}{\tau}\mathbb{E}_{(x,y)\sim p_{\text{pos}}}[f(x)^{\top}f(y)]$
$+\mathbb{E}_{x\sim p_{\text{data}}}\big{[}\log\mathbb{E}_{x^{-}\sim p_{\text{data}}}[e^{f(x^{-})^{\top}f(x)/\tau}]\big{]}.$ (11)
Hence, the contrastive learning signal can be decomposed into two components: the alignment term and the uniformity term. The alignment term minimizes the distance between embeddings of positive pairs, while the uniformity term encourages negative embeddings to be more uniformly distributed. In CARFT, the contrastive learning signal helps align the on-policy sampled CoT distribution with the annotated CoT. This effectively leverages the information contained in the annotated CoT and enhances training stability.